# A contact topological glossary for non-equilibrium thermodynamics

### v2

@@ 5 Slow global processes and partial order on Legendrians
Let $[\Lambda]$ be the family of all Legendrians in $(\Sigma,\xi)$ that can be obtained from a given properly embedded Legendrian $\Lambda$ by a compactly supported Legendrian isotopy. We say that a compactly supported Legendrian isotopy $\Lambda_{t}$, $t\in[0,1]$, of Legendrians in $[\Lambda]$ is non-negative if it admits a parameterization with a vector field $v_{t}$ satisfying $\alpha(v_{t})\geq 0$ (note that if this holds for one parameterization, then it holds for all the others). We write $\Lambda_{0}\preceq\Lambda_{1}$ if there exists a non-negative Legendrian isotopy from $\Lambda_{0}$ to $\Lambda_{1}$.
@@
The class $[\Lambda]$ is called orderable if the binary relation $\preceq$ is a partial order. A Reeb chord between two Legendrians is called non-trivial if it has positive length (as opposed to a point). The space of intensive variables $X$ below is assumed to be a connected manifold without boundary (either closed, i.e. compact, or open, i.e., non-compact).
@@
For any (open or closed) smooth connected manifold $X$, the class $[0_{X}]$ of the zero section $0_{X}$ in $J^{1}X=T^{*}X\times{\mathbb{R}}$ is orderable. If two distinct equilibrium Legendrians $\Lambda_{0},\Lambda_{1}\in[0_{X}]$ are related by the partial order, $\Lambda_{0}\preceq\Lambda_{1}$, there is no admissible global process taking $\Lambda_{1}$ to $\Lambda_{0}$.
@@
If two distinct Legendrians $\Lambda_{0},\Lambda_{1}\in[0_{X}]$ satisfy $\Lambda_{0}\preceq\Lambda_{1}$, there exists a non-trivial Reeb chord starting on $\Lambda_{0}$ and ending on $\Lambda_{1}$. In particular, if two equilibrium thermodynamic Legendrians from $[0_{X}]$ are related by an admissible global process, then there is an ultrafast process which starts at an equilibrium state of the first system and ends at an equilibrium state of the second system.
@@
For every two distinct Legendrians in $[0_{X}]$ there exists a non-trivial Reeb chord connecting them (in an unspecified order). If these are the equilibrium Legendrians of two thermodynamic systems, there exists an ultrafast process that starts at an equilibrium state of one system and ends at an equilibrium state of the other.
@@
In this case, Theorem 5.1 follows from [5, Proposition 5], [7, Theorem 1]. Theorem 5.2 can be deduced from the theory of generating functions and their spectral selectors [6, 7, 29] along the following lines. Assume without loss of generality that $\Lambda_{0}$ is the zero section. Let $c_{+}$ and $c_{-}$ be the spectral selectors corresponding to the fundamental class (i.e., the class of $X$) and the class of the point in the homology $H_{*}(X,\mathbb{Z}_{2})$ with the $Z_{2}$-coefficients, respectively. If $\Lambda_{0}\preceq\Lambda_{1}$, then $c_{+}(\Lambda_{1})\geq 0$. Since $\Lambda_{0}\neq\Lambda_{1}$, we have $c_{+}(\Lambda_{1})>0$. It follows from the spectrality property of the spectral selectors (see e.g. Proposition 2.4(i) in [29]) that there exists a chord of time-length $c_{+}(\lambda_{1})$ starting on $\Lambda_{0}$ and ending on $\Lambda_{1}$, as required. Theorem 5.3 follows from the fact that since $\Lambda_{0}\neq\Lambda_{1}$, at least one of $c_{-}(\Lambda_{1})$ or $c_{+}(\Lambda_{1})$ is distinct from 00.
A different argument which deduces Theorems 5.2 and 5.3 from the orderability of the class $[0_{X}]$ is discovered in a recent paper by Allais and Arlove [1], see Corollary 1.5 and Theorem 4.7 therein. Interestingly enough, this argument works on arbitrary contact manifolds, with $[0_{X}]$ being replaced by arbitrary orderable Legendrian isotopy class of a closed Legendrian submanifold. Our exposition is influenced by this paper.
@@ 6.1 An isochoric temperature and pressure jump
Consider the lattice gas consisting of $N$ particles occupying $V$ sites at temperature $T$ in the thermodynamic limit $V,N\to+\infty$ so that $V/N\to v$. Furthermore, we assume that the gas is diluted, $v\gg 1$. This is a model approximating the ideal gas – see [13, Section 1.3.2] for details.
@@
$(P+P_{back})v=T.$ (29)
Here, $P_{back}$ is a background pressure that we treat as a physical parameter of the system. If the molecules do not interact so that the internal energy vanishes, then the Hamiltonian in this model is linear, of the form (3)-(4) with $V_{int}(m)=0$. Then, a straightforward calculation following [13] shows that the equilibrium free energy $G$ in the ideal gas approximation is
$G=-T\ln v=T\ln((P+P_{back})/T).$
@@
Assume now that the temperature jumped from $T_{0}$ to $T_{1}>T_{0}$ and the background pressure simultaneously jumped from $P_{back}=0$ to $P_{back}=c>0$, while the volume is kept fixed. One can think that the gas is contained in a chamber with one of the walls being a piston of area $a$, and one instantly changes the temperature and simultaneously switches on a force $f$ applied to the piston and directed towards the gas so that $c=f/a$.
Given the parameters $T_{0}$, $T_{1}$ and $c$, there exist unique values of the volume $v$ and the initial pressure $P_{0}$ when such a jump corresponds to the ultrafast process, see Figure 1. This happens if, after the jump, the system lands in the equilibrium of the terminal system, so that
@@
In particular, if the background pressure jump $c\ll 1$ is small, and $T_{1}-T_{0}=O(1)$, the volume $v$ is large in accordance to the ideal gas approximation.
Figure 1: $T_{0}=1,T_{1}=5,c=2$
@@
When discussing an instantaneous isochoric process, as in the Stirling engine, temperature acts as a parameter, while pressure is a macroscopic variable. Introducing an artificial parameter, the background pressure, such that the total pressure is given by $P+P_{back}$, allows us to describe the simultaneous jump of two parameters: $T$ jumps from $T_{0}$ to $T_{1}$, and $P_{back}$ from 00 to $c$. More generally, this trick applies to processes in which intensive variables exhibit instantaneous jumps while the corresponding extensive variables remain unchanged, see Section 6.2 for another example. This not only streamlines the exposition, but, more importantly, gives rise to the connection between such processes and Reeb chords.
Let us explain now how such ultrafast process corresponds to a Reeb chord between two Legendrian submanifolds. We work in the reduced thermodynamic phase space $(z,p,q)$ where $z=-G,p=v,q=-P$. This notation is in accordance with the fact that volume $v$ is an extensive variable and the pressure $P$ is the intensive variable [^5] We hope that the difference between $p$ and $P$ will not cause confusion for the reader..
@@
$\Lambda(T,P_{back})=\{(z,p,q)\in{\mathbb{R}}^{3}\;:\;z=\phi_{T}(q-P_{back}),\;p=\phi^{\prime}_{T}(q-P_{back})\}\;,$ (33)
@@
which coincides with (30). Thus, the ultrafast process discussed above, corresponds exactly to the Reeb chord connecting $\Lambda_{0}$ with $\Lambda_{1}$. Furthermore, the orientation of the chord is given by $\partial/\partial z$ since $z_{1}-z_{0}>0$ by (32) and, as we recall, $z=-G$.
@@
The projections of the resulting Legendrians and the chord onto the $(q,z)$ plane (the front projection) are given on Fig. 3.
Figure 3: $T_{0}=1,T_{1}=5,c=2$
@@
It sounds likely that there is no non-negative Legendrian isotopy (that is, a slow global thermodynamic process) connecting $\Lambda_{0}$ and $\Lambda_{1}$. To make this statement rigorous, one has to find a way to deal with the lack of compactness in this example, cf. Remark 2.1 above. For instance, let us modify $\overline{\Lambda}_{1}$ below $\overline{\Lambda}_{0}$ so that it coincides with the line $\{z=-1,p=0\}$ outside a compact set. Twist now $J^{1}{\mathbb{R}}$ into $J^{1}S^{1}$ by taking the quotient by $(z,p,q)\mapsto(z,p,q+C)$ with a large constant $C$, so that $\overline{\Lambda}_{0}$ and $\overline{\Lambda}_{1}$ are modified to the Legendrian circles $\Lambda^{\prime}_{0}$ and $\Lambda^{\prime}_{1}$. One can show that these circles are incomparable with respect to the partial order. This can be proved with the help of Legendrian spectral invariants which are monotone with respect to non-negative Legendrian isotopies, see [7, Section 4.1].
@@
Here, we illustrate that Reeb chords can also correspond to a slow thermodynamic path, still on the ideal gas example. Suppose the temperature increases slowly with time $0\leq t\leq 1$, given by $T(t)=T_{0}+(T_{1}-T_{0})t$, and the external pressure also increases slowly $P_{ext}(t)=ct$, while the volume remains constant, $v(t)=v$. The Lagrangian projection of the equilibrium Legendrian at time $t$ onto the $(p,q)$-plane is given by
@@ 6.2 A temperature and magnetic field jump with constant magnetization
We start with a discussion based on [12]. Consider the Curie-Weiss magnet in the thermodynamic limit, see Example 3.3. As the external magnetic field $H$ is an intensive variable, and the magnetization $M$ is an extensive one, we use the notation $p=M\in{\mathbb{R}}$ and $q=H\in{\mathbb{R}}$. Recall (cf. also Remark 6.1) that the total magnetic field is given by $H+H_{back}$, where $H_{back}$ is the background magnetic field, playing a role of the parameter. Suppose that the temperature of the Curie-Weiss magnet jumped from $T_{0}$ to $T_{1}>T_{0}$, and that the background magnetic field jumped from 00 to $c\in{\mathbb{R}}$. The reduced Legendrian equilibrium submanifold $\Lambda(T,H_{back})$ is defined by the first two equations in (26). Let us set $\Lambda_{0}=\Lambda(T_{0},0)$, and let $\Lambda_{1}=\Lambda(T_{1},c)$.
@@
The Legendrian $\Lambda(T,H_{back})$ can be equivalently represented as
$z=\phi_{T}(q+H_{back}+bp)-\frac{bp^{2}}{2},~{}~{}p=\phi_{T}^{\prime}(q+H_{back}+bp).$ (38)
@@
Let now $g_{t}$, $t\in[0,1]$ be a compactly supported contact isotopy of $J^{1}{\mathbb{R}}$ satisfying (37). Fix sufficiently large positive numbers $R>r>0$, let $A_{R}$ be the cube $[-R,R]^{3}\subset{\mathbb{R}}^{3}=J^{1}{\mathbb{R}}$ and denote by $S_{R}$ the circle ${\mathbb{R}}/R{\mathbb{Z}}$ in the $Q$-variable. For a Legendrian submanifold $Y$ which coincides with $K$ outside $A_{r}$, we denote by $\widehat{Y}$ the Legendrian submanifold of $J^{1}S_{R}$ obtained by the following procedure (see Fig. 5). First, interpolate between $Y$ and the Legendrian submanifold
@@
inside the interior of $A_{R}$ without creating new chords from the zero section $L$ to the resulting submanifold $\widehat{Y}$. Then, glue together the points of this submanifold corresponding to $Q=-R$ and $Q=R$. This is possible simply because $L_{c}$ does not depend on $Q$.

### v3

@@ 5 Slow global processes and partial order on Legendrians
Let $\Lambda]$ be the family of all Legendrians in $(\Sigma,\xi)$ that can be obtained from a given properly embedded Legendrian $\Lambda$ by a compactly supported Legendrian isotopy. We say that a compactly supported Legendrian isotopy $\Lambda_{t}$, $t\in[0,1]$, of Legendrians in $\Lambda]$ is non-negative if it admits a parameterization with a vector field $v_{t}$ satisfying $\alpha(v_{t})\geq 0$ (note that if this holds for one parameterization, then it holds for all the others). We write $\Lambda_{0}\preceq\Lambda_{1}$ if there exists a non-negative Legendrian isotopy from $\Lambda_{0}$ to $\Lambda_{1}$.
@@
The class $\Lambda]$ is called orderable if the binary relation $\preceq$ is a partial order. A Reeb chord between two Legendrians is called non-trivial if it has positive length (as opposed to a point). The space of intensive variables $X$ below is assumed to be a connected manifold without boundary (either closed, i.e. compact, or open, i.e., non-compact).
@@
For any (open or closed) smooth connected manifold $X$, the class $0_{X}]$ of the zero section $0_{X}$ in $J^{1}X=T^{*}X\times{\mathbb{R}}$ is orderable. If two distinct equilibrium Legendrians $\Lambda_{0},\Lambda_{1}\in[0_{X}]$ are related by the partial order, $\Lambda_{0}\preceq\Lambda_{1}$, there is no admissible global process taking $\Lambda_{1}$ to $\Lambda_{0}$.
@@
If two distinct Legendrians $\Lambda_{0},\Lambda_{1}\in[0_{X}]$ satisfy $\Lambda_{0}\preceq\Lambda_{1}$, there exists a non-trivial Reeb chord starting on $\Lambda_{0}$ and ending on $\Lambda_{1}$. In particular, if two equilibrium thermodynamic Legendrians from $0_{X}]$ are related by an admissible global process, then there is a free-energy non-increasing ultrafast process which starts at an equilibrium state of the first system and ends at an equilibrium state of the second system.
@@
For every two distinct Legendrians in $0_{X}]$ there exists a non-trivial Reeb chord connecting them (in an unspecified order). If these are the equilibrium Legendrians of two thermodynamic systems, there exists an ultrafast process that starts at an equilibrium state of one system and ends at an equilibrium state of the other.
@@
In this case, Theorem 5.1 follows from [6, Proposition 5] and [8, Theorem 1]. Theorem 5.2 can be deduced from the theory of generating functions and their spectral selectors [7, 8, 38] along the following lines. Assume without loss of generality that $\Lambda_{0}$ is the zero section. Let $c_{+}$ and $c_{-}$ be the spectral selectors corresponding to the fundamental class (i.e., the class of $X$) and the class of the point in the homology $H_{*}(X,\mathbb{Z}_{2})$ with the $\mathbb{Z}_{2}$-coefficients, respectively. If $\Lambda_{0}\preceq\Lambda_{1}$, then $c_{+}(\Lambda_{1})\geq 0$. Since $\Lambda_{0}\neq\Lambda_{1}$, we have $c_{+}(\Lambda_{1})>0$. It follows from the spectrality property of the spectral selectors (see e.g. Proposition 2.4(i) in [38]) that there exists a chord of time-length $c_{+}(\lambda_{1})$ starting on $\Lambda_{0}$ and ending on $\Lambda_{1}$, as required. Theorem 5.3 follows from the fact that since $\Lambda_{0}\neq\Lambda_{1}$, at least one of $c_{-}(\Lambda_{1})$ or $c_{+}(\Lambda_{1})$ is distinct from 0.
A different argument which deduces Theorems 5.2 and 5.3 from the orderability of the class $0_{X}]$ is discovered in a recent paper by Allais and Arlove [1], see Corollary 1.5 and Theorem 4.7 therein. Interestingly enough, this argument works on arbitrary contact manifolds, with $0_{X}]$ being replaced by arbitrary orderable Legendrian isotopy class of a closed Legendrian submanifold. Our exposition is influenced by this paper.
@@ 6.1 An isochoric temperature and pressure jump
Consider the lattice gas consisting of $N$ particles occupying $V$ sites at temperature $T$ in the thermodynamic limit $V,N\to+\infty$ so that $V/N\to v$. Furthermore, we assume that the gas is diluted, $v\gg 1$. This is a model approximating the ideal gas – see [15, Section 1.3.2] for details.
@@
$(P+P_{\rm back})v=T.$ (33)
Here, $P_{\rm back}$ is a background pressure that we treat as a physical parameter of the system. If the molecules do not interact so that the internal energy vanishes, then the Hamiltonian in this model is linear, of the form (3)-(4) with $V_{\rm int}(m)=0$. Then, a straightforward calculation following [15] shows that the equilibrium free energy $G$ in the ideal gas approximation is
$G=-T\ln v=T\ln((P+P_{\rm back})/T).$
@@
Assume now that the temperature jumped from $T_{0}$ to $T_{1}>T_{0}$ and the background pressure simultaneously jumped from $P_{\rm back}=0$ to $P_{\rm back}=c>0$, while the volume is kept fixed. One can think that the gas is contained in a chamber with one of the walls being a piston of area $a$, and one instantly changes the temperature and simultaneously switches on a force $f$ applied to the piston and directed towards the gas so that $c=f/a$.
Given the parameters $T_{0}$, $T_{1}$ and $c$, there exist unique values of the volume $v$ and the initial pressure $P_{0}$ when such a jump corresponds to the ultrafast process. This happens if, after the jump, the system lands in the equilibrium of the terminal system, so that
@@
In particular, if the background pressure jump $c\ll 1$ is small, and $T_{1}-T_{0}=O(1)$, the volume $v$ is large in accordance to the ideal gas approximation. Recall that this ultrafast process corresponds to the Reeb chord connecting initial and terminal equilibrium Legendrians. In the $(v,P)$-plain this chord corresponds to the intersection point of the Lagrangian projections of these Legendrians. This is visualized on Figure 1.
Figure 1: Ultrafast process for $T_{0}=1,T_{1}=5,c=2$
@@
Our model ignores the Maxwell–Boltzmann distribution of particle velocities. A possible objection to this simplification is that the Maxwell–Boltzmann distribution depends on temperature, so a temperature jump would inevitably drive the system out of equilibrium. To address this, we tacitly assume that relaxation of the velocities to equilibrium is much faster than that of the coordinates. The physical feasibility of this assumption will be discussed elsewhere.
Remark 6.2.
When discussing an instantaneous isochoric process, as in the Stirling engine, temperature acts as a parameter, while pressure is a macroscopic variable. Introducing an artificial parameter, the background pressure, such that the total pressure is given by $P+P_{\rm back}$, allows us to describe the simultaneous jump of two parameters: $T$ jumps from $T_{0}$ to $T_{1}$, and $P_{\rm back}$ from 0 to $c$. More generally, this trick applies to processes in which intensive variables exhibit instantaneous jumps while the corresponding extensive variables remain unchanged, see Section 6.2 for another example. This not only streamlines the exposition, but, more importantly, gives rise to the connection between such processes and Reeb chords.
Let us explain now how such ultrafast process corresponds to a Reeb chord between two Legendrian submanifolds. We work in the reduced thermodynamic phase space $(z,p,q)$ where z=G,p=v,q=Pz=-G,p=v,q=-P. This notation is in accordance with the fact that volume $v$ is an extensive variable and the pressure $P$ is the intensive variable [^5] We hope that the difference between $p$ and $P$ will not cause confusion for the reader..
@@
$\Lambda(T,P_{\rm back})=\{(z,p,q)\in{\mathbb{R}}^{3}\;:\;z=\phi_{T}(q-P_{\rm back}),\;p=\phi^{\prime}_{T}(q-P_{\rm back})\}\;,$ (37)
@@
which coincides with (34). Thus, the ultrafast process discussed above, corresponds exactly to the Reeb chord connecting $\Lambda_{0}$ with $\Lambda_{1}$. Furthermore, the orientation of the chord is given by $\partial/\partial z$ since $z_{1}-z_{0}>0$ by (36) and, as we recall, z=Gz=-G.
@@
The projections of the resulting Legendrians and the chord onto the $(q,z)$ - plane (the front projection) are given on Fig. 3.
Figure 3: Reeb chord for $T_{0}=1,T_{1}=5,c=2$
@@
It sounds likely that there is no non-negative Legendrian isotopy (that is, a slow global thermodynamic process) connecting $\Lambda_{0}$ and $\Lambda_{1}$. To make this statement rigorous, one has to find a way to deal with the lack of compactness in this example, cf. Remark 2.2 above. For instance, let us modify $\overline{\Lambda}_{1}$ below $\overline{\Lambda}_{0}$ so that it coincides with the line $\{z=-1,p=0\}$ outside a compact set. Twist now $J^{1}{\mathbb{R}}$ into $J^{1}S^{1}$ by taking the quotient by $(z,p,q)\mapsto(z,p,q+C)$ with a large constant $C$, so that $\overline{\Lambda}_{0}$ and $\overline{\Lambda}_{1}$ are modified to the Legendrian circles $\Lambda^{\prime}_{0}$ and $\Lambda^{\prime}_{1}$. One can show that these circles are incomparable with respect to the partial order. This can be proved with the help of Legendrian spectral invariants which are monotone with respect to non-negative Legendrian isotopies, see [8, Section 4.1].
@@
Here, we illustrate that Reeb chords can also correspond to a slow thermodynamic path, still on the ideal gas example. Suppose the temperature increases slowly with time $0\leq t\leq 1$, given by $T(t)=T_{0}+(T_{1}-T_{0})t$, and the external pressure also increases slowly $P_{\rm ext}(t)=ct$, while the volume remains constant, $v(t)=v$. The Lagrangian projection of the equilibrium Legendrian at time $t$ onto the $(p,q)$-plane is given by
@@ 6.2 A temperature and magnetic field jump with constant magnetization
We start with a discussion based on [13]. Consider the Curie-Weiss magnet in the thermodynamic limit, see Example 3.3. As the external magnetic field $H$ is an intensive variable, and the magnetization $M$ is an extensive one, we use the notation $p=M\in{\mathbb{R}}$ and $q=H\in{\mathbb{R}}$. Recall (cf. also Remark 6.2) that the total magnetic field is given by $H+H_{\rm back}$, where $H_{\rm back}$ is the background magnetic field, playing a role of the parameter. Suppose that the temperature of the Curie-Weiss magnet jumped from $T_{0}$ to $T_{1}>T_{0}$, and that the background magnetic field jumped from 0 to $c\in{\mathbb{R}}$. The reduced Legendrian equilibrium submanifold $\Lambda(T,H_{\rm back})$ is defined by the first two equations in (26). Let us set $\Lambda_{0}=\Lambda(T_{0},0)$, and let $\Lambda_{1}=\Lambda(T_{1},c)$.
@@
The Legendrian $\Lambda(T,H_{\rm back})$ can be equivalently represented as
$z=\phi_{T}(q+H_{\rm back}+bp)-\frac{bp^{2}}{2},~~p=\phi_{T}^{\prime}(q+H_{\rm back}+bp).$ (42)
@@
Let now $g_{t}$, $t\in[0,1]$ be a compactly supported contact isotopy of $J^{1}{\mathbb{R}}$ satisfying (41). Fix sufficiently large positive numbers $R>r>0$, let $A_{R}$ be the cube $R,R]^{3}\subset{\mathbb{R}}^{3}=J^{1}{\mathbb{R}}$ and denote by $S_{R}$ the circle ${\mathbb{R}}/R{\mathbb{Z}}$ in the $Q$-variable. For a Legendrian submanifold $Y$ which coincides with $K$ outside $A_{r}$, we denote by $\widehat{Y}$ the Legendrian submanifold of $J^{1}S_{R}$ obtained by the following procedure (see Fig. 5). First, interpolate between $Y$ and the Legendrian submanifold
@@
inside the interior of $A_{R}$ without creating new chords from the zero section $L$ to the resulting submanifold $\widehat{Y}$. Then, glue together the points of this submanifold corresponding to Q=RQ=-R and $Q=R$. This is possible simply because $L_{c}$ does not depend on $Q$.
@@
6.3 An ultrafast (de)magnetization for the Curie-Weiss model
Consider the Curie-Weiss magnet at temperature $T$. Assume that the background magnetic field vanishes. Put $\psi(x)=T\ln\left(2\cosh(x/T)\right)$. Recall from Example 3.3 that in the equilibrium
$z=\psi(q+bp)-\frac{bp^{2}}{2}\;,$ (47)
$p=\psi^{\prime}(q+bp)\;,$ (48)
and
q=bp+Tu(p),q=-bp+Tu(p)\;, (49)
with
$u(p)=\frac{1}{2}\ln\frac{1+p}{1-p}\;.$
We view $z$ and $p$ as functions of $b$, with $T$ and $q$ being fixed. Differentiating (47) by $b$ and substituting (48) we get that
$\frac{dz}{db}=p^{2}/2\;.$
Furthermore, by (49)
$T^{-1}b=\frac{u(p)-T^{-1}q}{p}\;.$
Differentiating by $p$ we get
$T^{-1}\frac{db}{dp}=\frac{\left(pu^{\prime}(p)-u(p)\right)+T^{-1}q}{p^{2}}\;.$
The expression in brackets in the nominator is the $y$-intercept of the function $u$ taken with the opposite sign. Since for $p>0$ the function $u$ is strictly convex with $u(0)=0$, its intercept is negative. It follows that for $p>0$, $dz/db>0$ and $db/dp>0$ (and hence $dp/db>0$).
In particular, if as the result of certain ultrafast process, the exchange parameter (called sometimes coupling) $b$ is decreasing, the free energy increases and magnetization drops; and if $b$ is increasing, the free energy decreases and the magnetization is increasing. In some magnetic systems the manipulation of the parameter $b$ can be achieved by ultrashort laser pulses applied to the system ( see [29, 30]; we note that none of these papers use mean-field models; we introduce them here as an approximation.)
Data availability statement. This paper does not use any data.
Conflict of interest statement. On behalf of all authors, the corresponding author states that there is no conflict of interest.